# An Effective Positivstellensatz over the Rational Numbers for Finite Semialgebraic Sets

### v1

@@ Abstract
We study the problem of representing multivariate polynomials with rational coefficients, which are nonnegative and strictly positive on finite semialgebraic sets, using rational sums of squares.
We focus on the case of finite semialgebraic sets $S$ defined by equality constraints, generating a zero-dimensional ideal $I$, and by nonnegative sign constraints.
[…]
@@
Finally, we analyze the bit complexity of the Sums-of-Squares representations for polynomials with coefficients in $\mathbb{Q}$, in the case the ideal is radical. The bitsize bounds are quadratic or cubic in the Bezout bound, and linear in the regularity, generalizing and improving previous results obtained for special zero dimensional ideals.
@@ 1 Introdution
In this paper, we will consider finite semialgebraic sets, focusing in particular on the existence of exact effective SoS representations. Instead of working with the specific field $\mathbb{Q}$ of rational numbers, we study polynomials with coefficients in a subfield $K$ of $\mathbb{R}$, that are strictly positive (or nonnegative) on a finite basic closed semialgebraic set $S\subset\mathbb{R}^{n}$ defined by polynomials with coefficients in $K$. We aim at analyzing the smallest field in which we can compute a SoS representation and at bounding the degree and the height (when $K=\mathbb{Q}$) of the coefficients of this SoS representation.
The precise problems we consider here are the following. Given a field $K\subset\mathbb{R}$, and
@@
two sets of polynomials in $n$ variables with coefficients in $K$, where the ideal $I=(\mathbf{h})\subset K[\mathbf{x}]$ generated by $\mathbf{h}$ is zero-dimensional, which then define the finite (basic closed) semialgebraic set
$S=S(\mathbf{g},\mathbf{h})=\{\,\xi\in\mathbb{R}^{n}\colon g_{i}(\xi)\geq 0\textup{ for }1\leq i\leq r\textup{ and }h_{j}(\xi)=0\textup{ for }1\leq j\leq s\,\}\ \subset\mathbb{R}^{n}.$ (2)
@@
    with $\omega_{i,k}\in K_{\geq 0}$ and $q_{i,k},p_{j}\in K[\mathbf{x}]$, and what are the degrees of $q_{i,k},p_{j}$?
@@
is a finite set, or equivalently, when the quotient ring $K[\mathbf{x}]/I$ is a finite dimensional $K$-vector space. In the sequel we will denote by $B$ a monomial basis of $K[\mathbf{x}]/I$, and by $\langle B\rangle_{K}$, the vector space spanned by $B$ in $K[\mathbf{x}]$. We also recall that $\mathbf{h}$ is a graded basis of the ideal $I$ when for all $p\in I$, there exist $p_{j}\in K[\mathbf{x}]$ with $\deg(p_{j})\leq\deg(p)-\deg(h_{j})$, $1\leq j\leq s$, such that $p=\displaystyle\sum_{j=1}^{s}p_{j}\,h_{j}.$
@@
Theorem B.
Let $K$, $I$ and $S$ be as in A, and assume that $I$ is a radical ideal. Then for $f\in K[\mathbf{x}]$, $f\geq 0$ on $S$ if and only if $f$ admits a SoS representation as in (3).
@@
If $f>0$ on $S$ (resp. $f\geq 0$ on $S$), then the following holds for a SoS representation as in (3) of $f$:
@@
If in addition, $\mathbf{h}=\{h_{1},\dots,h_{s}\}$ is a graded basis of $I$, then $\deg(q_{i,j})\leq nd$ (resp. $\deg(q_{i,j})\leq(n+1)d$), $p_{j}\in\mathbb{Q}[\mathbf{x}]$ with $\deg(p_{j})\leq\widehat{d}-\deg(h_{j})$ and height bounded by $\widetilde{\mathcal{O}}(d^{2n})\,\delta\,\tau+\widetilde{\mathcal{O}}({\widehat{d}}^{\,n})\tau$, for $1\leq j\leq s$, where $\widehat{d}:=2(d+\deg(B))+1$.
@@
The bounds in C involve Bézout type bounds derived from [KPS01] for the height of the Chow form of $V_{\mathbb{C}}$. These bounds could be refined using mixed-volume bounds as in [EMT20] or [MS18], at the cost of a genericity assumption: we won’t follow this path hereafter.
@@ 1.1 Related works and Putinar-type representations
The quadratic module generated by the polynomials $\mathbf{g}=\{g_{1},\ldots,g_{r}\}$ and ±$\pm\mathbf{h}=\{\pm h_{1},\ldots,\pm h_{s}\}$ is
@@
Putinar’s Positivstellensatz [Put93] states that any real polynomial that is is strictly positive on a compact semialgebraic set $S$ can be represented in the quadratic module $Q_{\mathbb{R}}$, under the assumption that $Q_{\mathbb{R}}$ is Archimedean, i.e. that there exists $g\in Q_{\mathbb{R}}$ such that $\{\xi\in\mathbb{R}:g(\xi)\geq 0\}$ is compact. While Schmüdgen’s Positivstellensatz (closely related to Putinar’s one) is known to hold over any Archimedean field $K\subset\mathbb{R}$, see e.g. [Sch02], it is not known if the same is true for Putinar’s Positivstellensatz, because it is uncertain if $Q_{\mathbb{R}}$ Archimedean implies that $Q_{K}$ is Archimedean.
@@
In our setting, $Q_{K}$ satisfies the Archimedean condition by taking $g=-\sum h_{j}^{2}\in Q_{K}$, but none of the conditions above are necessarily satisfied.
@@
[…]
We now discuss the minimality of our assumption $(I:f)+(f)=(1)$. For zero-dimensional ideals, as already mentioned, $Q_{\mathbb{R}}$ is an Archimedean quadratic module (see also [Mar08, Cor. 7.4.4]), and therefore, by [Mar08, 9.1.2]
@@
The degrees of Putinar-type representations for strictly positive polynomials on general $S$ has been investigated in [NS07, MSED21] and more recently in [BM22, BMP25], respectively with exponential bounds and polynomial bounds depending on the minimum $f^{*}$ of $f$ on $S$: the closer $f$ is having a zero on $S$, the bigger is the degree needed for the representation.
@@
Regarding the optimality of these bounds, it is shown in [Gri01], that the degree of a SoS certificate for deciding the emptyness of a basic semialgebraic set is at least $c\,d^{n}$ for some $c\in\mathbb{R}_{>0}$. As discussed in Section 1.2, the bounds on the degree in C are tight in the case of binary polynomials $h_{i}=x_{i}^{2}-x_{i}$ and linear constraints $g_{i}$ (see also [KLM15]).
@@
Obtaining good bitsize bounds in real algebra is a (mostly open) research challenge. For instance, for the (more general) problem of certifying the insatisfyiability of sign conditions using the Krivine-Stengle Positvstellensatz, the only known bitsize bound, provided in the recent seminal work [LPR20], reads as:
@@
Determining lower bounds for the bitsize (and also the degrees) of SoS representations is a mostly open challenge with important applications. For instance, in Theoretical Computer Science, there is interest in studying the question of certifying emptiness of a given (finite) semialgebraic set $S$, by providing a SoS representation for $-1$, and the bitsize of such SoS representation governs the complexity of the problem. It has also been observed in [O’D17, RW17] that the bitsize of Positivstellensätze representations can grow exponentially in the number of variables. Lower bounds on the bitsize for low degree representations are also provided in these papers, showing in a particular example where $d=2$ that any degree $2$ SoS representations of a nonnegative polynomial must have bitsize in $\Theta(d^{n})$. Unfortunately, this lower bound does not apply in our case, since the given generators of the ideal of this example do not form a graded basis. We also notice that in [RW17, Th. 1], the existence of a SoS representation over $\mathbb{Q}$ is assumed, whereas Theorems A and B provide existential results for such representations. Finally, we mention that the tradeoff between bitsizes and degrees of representations in the Positivstellensätze has been investigated in [AH19, Hak21].
@@ 1.2 Applications to discrete polynomial optimization
An important area of applications of SoS representation is Polynomial Optimization, which relies on the celebrated moment/SoS hierarchy [Las01, Par00], that we recall briefly.
To define this hierarchy, we need the truncated quadratic module generated, in degree $\leq d$, by the polynomials $\mathbf{g}$ and ±$\pm\mathbf{h}$:
@@
The finite convergence of different families of convex relaxations for $0/1$ programs has been studied since a long time, see e.g. [SA90, LS91]. The comparison of these techniques with the more recent moment/SoS hierarchy has been performed in [Lau03], demonstrating the strength of this new approach. In [Lau06], the same degree bound of Corollary 1.1 is provided for the finite convergence of the Moment hierarchy (i.e. $f^{(\ell)}$) and the SoS hierarchy (i.e. $f_{(\ell)}$), without deducing degree bounds for the SoS representation (when it exists). Similarly, sparse-type bounds $12\ell\leq n+\deg(f)-$ 1 for the order $\ell$ of moment-like relaxations in the quotient algebra where $h_{i}=x_{i}^{2}-1$ and $V_{\mathbb{R}}=\{-1,1\}\textdegree^{n}$ have been obtained in [STKI17] (see also [FSP15]), without bounding the degree of SoS representations.
The bound of Corollary 1.1 is tight in general, see for instance [KLM15] and references therein: for binary optimization problems with $h_{i}=x_{i}^{2}-x_{i}$, finite convergence always happens at order $r=n$ when there are no inequalities, and at order $r=n+1$ when the $g_{i}$’s are linear. Applying Corollary 1.1 (and assuming that $\deg(f)\leq n$, which can be done by reducing $f$ modulo $\mathbf{h}$) we get the same bounds, and we furthermore show that a rational certificate of nonnegativity for $f-f^{*}$ exists. The cases when $f_{(n-1)}<f^{*}$ (resp. $f_{(n)}<f^{*}$) are identified and characterized in [KLM15], showing that SoS representations may not exists in degree $<2\delta$ (resp. $<2\delta+\max_{i}\deg(g_{i})$) and that the degree bounds of A (and Corollary 1.1) are tight.
@@ 1.3 Proof strategy and structure of the manuscript
The proof of A is developed in Sections 2 and 3. We first consider in Section 2.2 the case when the ideal $I$ is radical and $f$ is strictly positive on the whole $V_{\mathbb{R}}(I)$, obtaining in Theorem 2.1. This is a refined version of A in the radical case, with an additional non-vanishing condition that will allow us to lift our construction to a non-radical ideal $I$ in Section 2.3. Then, we consider in Section 2.4 the case when $f>0$ on an arbitrary finite basic closed semialgebraic set $S\subset\mathbb{R}^{n}$ as in (2), by perturbing the original polynomial $f$, which is positive on $S$, to an appropriate polynomial positive on all $V_{\mathbb{R}}(I)$. Finally, we study in Section 3 the case where $f\geq 0$ on $S$ with $1\in(I:f)+(f)$. In this case we show that we can replace our original polynomial $f$, which is nonnegative on $S$, by another related polynomial which is strictly positive on $S$, and conclude by applying the previous results.
@@
Finally Section 7 presents an algorithm to obtained sums of squares representations of the kind previously studied, and some examples that illustrate our constructions.

### v2

@@ Abstract
We study the problem of representing multivariate polynomials with rational coefficients, which are nonnegative and strictly positive on finite semialgebraic sets, using rational sums of squares. We focus on the case of finite semialgebraic sets $S$ defined by equality constraints, generating a zero-dimensional ideal $I$, and by nonnegative sign constraints.
First, we obtain existential results. We prove that a strictly positive polynomial $f$ with coefficients in a subfield $K$ of $\mathbb{R}$ has a representation in terms of weighted Sums-of-Squares with coefficients in this field, even if the ideal $I$ is not radical. We generalize this result to the case where $f$ is nonnegative on $S$ and $(f)+(I:f)=1$. We deduce that nonnegative polynomials with coefficients in $K$ can be represented in terms of Sum-of-Squares of polynomials with coefficients in $K$, when the ideal is radical.
@@
Finally, we analyze the bit complexity of the Sums-of-Squares representations for polynomials with coefficients in $\mathbb{Q}$, in the case the ideal is radical. The bitsize bounds are linear in the bitsize of the input polynomials, quadratic or cubic in the Bezout bound, and linear in the regularity, generalizing and improving previous results obtained for special zero dimensional ideals.
@@ 1 Introdution
In this paper, we will consider finite semialgebraic sets, focusing in particular on the existence of exact effective SoS representations. Instead of working with the specific field $\mathbb{Q}$ of rational numbers, we study polynomials with coefficients in a subfield $K$ of $\mathbb{R}$, that are strictly positive (or nonnegative) on a finite basic closed semialgebraic set $S\subset\mathbb{R}^{n}$ defined by polynomials with coefficients in $K$. We aim at analyzing the smallest field in which we can compute an SoS representation and at bounding the degree and the height (when $K=\mathbb{Q}$) of the coefficients of this SoS representation. In the case $K=\mathbb{R}$, these representations are well-studied, see [Par02] and Section 1.1. The main technical difficulty in generalizing these results to subfields of $\mathbb{R}$ is to ensure a certain strict feasibility condition for the matrices associated with the SoS representation: we refer to Section 2.2 for more details.
The precise problems we consider here are the following. Given a field $K\subset\mathbb{R}$, let
@@
be two sets of polynomials in $n$ variables with coefficients in $K$, where the ideal $I=(\mathbf{h})\subset K[\mathbf{x}]$ generated by $\mathbf{h}$ is zero-dimensional, i.e. with finitely many complex solutions. The polynomials $\mathbf{g}$ and $\mathbf{h}$ then define the finite (basic, closed) semialgebraic set
[…]
@@
    with $\omega_{i,k}\in K_{\geq 0}{:=\{\,\omega\in K\colon\omega\geq 0\,\}}$ and $q_{i,k},p_{j}\in K[\mathbf{x}]$, and what are the degrees of $q_{i,k},p_{j}$?
@@
is a finite set, or equivalently, when the quotient ring $K[\mathbf{x}]/I$ is a finite dimensional $K$-vector space. In the sequel we will denote by $B$ a monomial basis of $K[\mathbf{x}]/I$, and by $\langle B\rangle_{K}$, the vector space spanned by $B$ in $K[\mathbf{x}]$. We also recall that $\mathbf{h}$ is a graded basis of the ideal $I$ when for all $p\in I$, there exist $p_{j}\in K[\mathbf{x}]$ with $\deg(p_{j})\leq\deg(p)-\deg(h_{j})$, $1\leq j\leq s$, such that $p=\displaystyle\sum_{j=1}^{s}p_{j}\,h_{j}$. Finally, for $f\in K[\mathbf{x}]$ we denote $(I\colon f)=\{\,h\in K[\mathbf{x}]\colon fh\in I\,\}$ the colon or quotient ideal.
@@
Corollary B.
Let $K$, II and $S$ be as in A, and assume that $I$ is a radical ideal. Then for $f\in K[\mathbf{x}]$, $f\geq 0$ on $S$ if and only if $f$ admits an SoS representation as in (3).
@@
If $f>0$ on $S$ (resp. $f\geq 0$ on $S$), then the following holds for an SoS representation as in (3) of $f$:
@@
If in addition, $\mathbf{h}=\{h_{1},\dots,h_{s}\}$ is a graded basis of $I$, then $\deg(q_{i,j})\leq nd$ (resp. $\deg(q_{i,j})\leq(n+1)d$), and $p_{j}\in\mathbb{Q}[\mathbf{x}]$ with degrees bounded by $\widehat{d}-\deg(h_{j})$ and heights bounded by $\widetilde{\mathcal{O}}(d^{2n})\,\delta\,\tau+\widetilde{\mathcal{O}}({\widehat{d}}^{\,n})\tau$, for $1\leq j\leq s$, where $\widehat{d}:=2(d+\deg(B))+1$.
@@
The bounds in C involve Bézout type bounds derived from [KPS01] for the height of the Chow form of $V_{\mathbb{C}}$. These bounds could be refined using mixed-volume bounds as in [MS18] or [EMT20], at the cost of a genericity assumption: we won’t follow this path hereafter.
@@ 1.1 Related works and Putinar-type representations
The quadratic module generated by the polynomials $\mathbf{g}=\{g_{1},\ldots,g_{r}\}$ and $\pm\mathbf{h}=\{\pm h_{1},\ldots,\pm h_{s}\}$ is
@@
Putinar’s Positivstellensatz [Put93] states that any real polynomial that is strictly positive on a compact semialgebraic set $S$ can be represented in the quadratic module $Q_{\mathbb{R}}$, under the assumption that $Q_{\mathbb{R}}$ is Archimedean, i.e. that there exists $g\in Q_{\mathbb{R}}$ such that $\{\xi\in\mathbb{R}:g(\xi)\geq 0\}$ is compact. While Schmüdgen’s Positivstellensatz (closely related to Putinar’s one) is known to hold over any Archimedean field $K\subset\mathbb{R}$, see e.g. [Sch02], it is not known if the same is true for Putinar’s Positivstellensatz, because it is uncertain if $Q_{\mathbb{R}}$ Archimedean implies that $Q_{K}$ is Archimedean.
@@
In our setting, $Q_{K}$ satisfies the Archimedean condition by taking $g=-\sum h_{j}^{2}\in Q_{K}$, but none of the conditions above are obviously satisfied: it is an open question whether the Archimedean condition over $\mathbb{Q}$ implies the condition in [Pow11], and the unisolvent property for all polynomials of bounded degree cannot occur if the semialgebraic set is zero-dimensional.
@@
As shown in Example 3.1, not all nonnegative polynomials on $S$ have a representation in $Q_{K}$. In the case of $K=\mathbb{R}$, Theorem A can be viewed as an extension of [Par02], which holds for nonnegative polynomials on finite semialgebraic sets $S$ defined by a radical zero-dimensional ideal $I$, to an arbitrary zero-dimensional ideal $I$ for $f$ satisfying $(I:f)+(f)=(1)$. Moreover, B can be seen as a full extension of the same paper’s result to other fields $K\subset\mathbb{R}$, and in particular to the field $\mathbb{Q}$.
Another related work over $\mathbb{R}$ is [HQ25]. In this work, it is assumed that the first $n$ equations $h_{1},\ldots,h_{n}$ have no root at infinity so that they form a graded basis (see [HQ25, Th. 3]). It is also assumed that the zeros of $I$ where $f-f_{*}$ vanishes are simple, which implies that $(I\colon(f-f_{*}))+(f-f_{*})=(1)$. The degree bounds on the SoS representation (see [HQ25, Th. 1]) depend on $\mathfrak{n}=\sum_{i=1}^{n}\deg(h_{i})-n$, which is the minimal degree of a basis of the quotient by $(h_{1},\ldots,h_{n})$. A generalizes the degree bounds of [HQ25] by providing degree bounds for any subfield $K\subset\mathbb{R}$ in terms of the minimal degree of a basis $B$ of the quotient by $(h_{1},\ldots h_{s})$, which can be substantially smaller than $\mathfrak{n}$, and by allowing a weaker regularity condition. See e.g. Example 7.3 for an example where A applies, but [HQ25, Th. 1] does not.
We now discuss the minimality of our assumption $(I:f)+(f)=(1)$ in general. For zero-dimensional ideals, as already mentioned, $Q_{\mathbb{R}}$ is an Archimedean quadratic module (see also [Mar08, Cor. 7.4.4]), and therefore, by [Mar08, 9.1.2]
@@
The degrees of Putinar-type representations for strictly positive polynomials on general $S$ have been investigated in [NS07, MSED21] and more recently in [BM23, BMP25], respectively with exponential bounds and polynomial bounds depending on the minimum $f^{*}$ of $f$ on $S$: the closer $f$ is having a zero on $S$, the bigger is the degree needed for the representation. These bounds have also been generalized to the matrix case, see [Hua25].
@@
Regarding the optimality of these bounds, it is shown in [Gri01], that the degree of an SoS certificate for deciding the emptyness of a basic semialgebraic set is at least $c\,d^{n}$ for some $c\in\mathbb{R}_{>0}$. As discussed in Section 1.2, the bounds on the degree in C are tight in the case of binary polynomials $h_{i}=x_{i}^{2}-x_{i}$ and linear constraints $g_{i}$ (see also [KLM15]).
@@
Obtaining good bitsize bounds in real algebra is a (mostly open) research challenge. For instance, for the (more general) problem of certifying the insatisfiability of sign conditions using the Krivine-Stengle Positvstellensatz, the only known bitsize bound, provided in the recent seminal work [LPR20], reads as:
@@
Determining lower bounds for the bitsize (and also the degrees) of SoS representations is a mostly open challenge with important applications. For instance, in Theoretical Computer Science, there is interest in studying the question of certifying emptiness of a given (finite) semialgebraic set $S$, by providing an SoS representation for 1-1, and the bitsize of such SoS representation governs the complexity of the problem. It has also been observed in [O’D17, RW17] that the bitsize of Positivstellensätze representations can grow exponentially in the number of variables. Lower bounds on the bitsize for low degree representations are also provided in these papers, showing in a particular example where $d=2$ that any degree $2$ SoS representation of a nonnegative polynomial must have bitsize in $\Theta(d^{n})$. Unfortunately, this lower bound does not apply in our case, since the given generators of the ideal of this example do not form a graded basis. We also notice that in [RW17, Th. 1], the existence of an SoS representation over $\mathbb{Q}$ is assumed, whereas Theorems A and B provide existential results for such representations. Finally, we mention that the tradeoff between bitsizes and degrees of representations in the Positivstellensätze has been investigated in [AH19, Hak21].
@@ 1.2 Applications to discrete polynomial optimization
An important area of applications of SoS representations is Polynomial Optimization, which relies on the celebrated moment/SoS hierarchy [Las01, Par00], that we recall briefly.
To define this hierarchy, we need the truncated quadratic module generated, in degree $\leq d$, by the polynomials $\mathbf{g}$ and $\pm\mathbf{h}$:
@@
The finite convergence of different families of convex relaxations for $0/1$ programs has been studied since a long time, see e.g. [SA90, LS91]. The comparison of these techniques with the more recent moment/SoS hierarchy has been performed in [Lau03], demonstrating the strength of this new approach. In [Lau06], the same degree bound of Corollary 1.1 is provided for the finite convergence of the Moment hierarchy (i.e. $f^{(\ell)}$) and the SoS hierarchy (i.e. $f_{(\ell)}$), without deducing degree bounds for the SoS representation (when it exists). Similarly, sparse-type bounds $2\ell\leq n+\deg(f)-1$ for the order $\ell$ of moment-like relaxations in the quotient algebra where $h_{i}=x_{i}^{2}-1$ and $V_{\mathbb{R}}=\{-1,1\}\textdegree^{n}$ have been obtained in [STKI17] (see also [FSP15]), without bounding the degree of SoS representations.
The bound of Corollary 1.1 is tight in general, see for instance [KLM15] and references therein: for binary optimization problems with $h_{i}=x_{i}^{2}-x_{i}$, finite convergence always happens at order $r=n$ when there are no inequalities, and at order $r=n+1$ when the $g_{i}$’s are linear. Applying Corollary 1.1 (and assuming that $\deg(f)\leq n$, which can be done by reducing $f$ modulo $\mathbf{h}$) we get the same bounds, and we furthermore show that a rational certificate of nonnegativity for $f-f^{*}$ exists. The cases when $f_{(n-1)}<f^{*}$ (resp. $f_{(n)}<f^{*}$) are identified and characterized in [KLM15], showing that SoS representations may not exists in degree $2\delta$ (resp. $2\delta+\max_{i}\deg(g_{i})$) and that the degree bounds of A (and Corollary 1.1) are tight.
@@ 1.3 Proof strategy and structure of the manuscript
The proof of A is developed in Sections 2 and 3. We first consider in Section 2.2 the case when the ideal $I$ is radical and $f$ is strictly positive on the whole $V_{\mathbb{R}}(I)$, obtaining Theorem 2.1. This is a refined version of A in the radical case, with an additional non-vanishing condition that will allow us to lift our construction to a non-radical ideal $I$ in Section 2.3. Then, we consider in Section 2.4 the case when $f>0$ on an arbitrary finite basic closed semialgebraic set $S\subset\mathbb{R}^{n}$ as in (2), by perturbing the original polynomial $f$, which is positive on $S$, to an appropriate polynomial positive on all $V_{\mathbb{R}}(I)$. Finally, we study in Section 3 the case where $f\geq 0$ on $S$ with $1\in(I:f)+(f)$. In this case we show that we can replace our original polynomial $f$, which is nonnegative on $S$, by another related polynomial which is strictly positive on $S$, and conclude by applying the previous results.
@@
Finally Section 7 presents an algorithm to obtain SoS representations of the kind previously studied, and some examples that illustrate our constructions.